# Utility-Diversity Aware Online Batch Selection for LLM Supervised Fine-tuning

### v1

@@ 1 Introduction
The rapid progress of Large Language Models (LLMs) has reshaped natural language processing and enabled impressive generalization across a wide range of domains (Achiam et al., 2023; Brown et al., 2020; Liu et al., 2024; Touvron et al., 2023; Bai et al., 2023). To further improve their performance in specialized areas, Supervised Fine-tuning (SFT) has emerged as a typical post-training paradigm. However, training on a full dataset is not always advantageous. It entails substantial computational cost and has been shown to be less effective than using a small amount of carefully curated data (Albalak et al., 2024; Xia et al., 2024; Zhou et al., 2023). These challenges highlight the importance of principled data selection strategies that can improve both efficiency and effectiveness in SFT.
@@
In this work, we address this challenge by focusing on online batch selection, a paradigm that dynamically evaluates sample value and performs filtering during training (Wang et al., 2024; Loshchilov & Hutter, 2015; Katharopoulos & Fleuret, 2018; Mindermann et al., 2022). Instead of using all samples in a batch, the model assesses the importance of each sample as it is encountered and selects only a subset to participate in parameter updates. This approach enables the selection process to adapt in real time to the model’s current state and learning trajectory, improving both training efficiency and effectiveness.
@@
However, despite promising progress, current online batch selection methods still face several issues that hinder their practical deployment. These methods typically rely solely on data utility, e.g., picking sample subsets with high loss (Loshchilov & Hutter, 2015; Jiang et al., 2019) or gradient magnitude (Katharopoulos & Fleuret, 2018). While such heuristics enable capturing the critical samples to reduce the current training loss, they evaluate sample value from a limited perspective. In practice, effective selection also requires considerations of intra-sample diversity, to reflect the richness of information with fewer repetitive phrases within each training instance, and inter-sample diversity to suppress redundancy across examples by avoiding repeated training on near-duplicate content (Lee et al., 2021; Tirumala et al., 2023). Popular literature (Loshchilov & Hutter, 2015; Katharopoulos & Fleuret, 2018; Mindermann et al., 2022) prioritizes data utility and rarely examines the selection criteria through the lens of diversity.
@@
Another primary issue involves both external dependencies and computational overhead. Some approaches (Mindermann et al., 2022; Wang et al., 2024) rely on a held-out validation set, which may be unavailable since we hardly know the exact distribution of test dataset. Other methods (Mindermann et al., 2022; Deng et al., 2023) use a reference model that is also shown to be impractical in the real world (Kaddour et al., 2023). Besides these external dependencies, several selection algorithms (Katharopoulos & Fleuret, 2018; Wang et al., 2024; Mindermann et al., 2022) introduce substantial computational costs that even exceed the expense of full-dataset training, raising efficiency concerns.
@@
Leveraging logits as basis for online scoring. To operationalize these desiderata, a key challenge lies in how to efficiently and reliably score candidate samples during training. In online batch selection, each candidate sample must be evaluated under the current model to quantify its contribution to learning. This evaluation typically requires at least a forward pass, as it can fully capture how the model presently interprets the sample and thus provide an accurate basis for selection (Huang et al., 2023; Shorinwa et al., 2025). To achieve this efficiently (D3) without external resources (D2), we must avoid expensive gradient computations for each candidate sample and leverage intrinsic signals already available during forward propagation. These considerations collectively justify leveraging the model’s output logits, which are naturally produced during the forward pass and encode rich information about both sample utility and diversity (Qiu & Miikkulainen, 2024; Geng et al., 2023). Specifically, we compute an intra-sample importance score based on the nuclear norm of the logits, which captures the utility and intra-sample diversity of a sample, and an inter-sample importance score using low-dimensional similarity matching against previously selected samples to quantify inter-sample diversity (D1). Integrating these scores allows the training process to balance exploitation of high-utility samples with exploration of less-visited regions in the data distribution, thereby improving the overall effectiveness of online batch selection.
@@ 3.1 Intra-sample Importance Score via Nuclear Norm
(Horn & Johnson, 2012) For any matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$, the following inequality holds:
@@
Larger Nuclear Norm indicates higher Optimization Utility.
Traditional notions of sample difficulty—such as maximum loss (Loshchilov & Hutter, 2015) or maximum gradient (Katharopoulos & Fleuret, 2018)—often misalign with actual training dynamics (Wang et al., 2024). Instead, we characterize sample importance through its potential contribution to loss reduction, termed its optimization utility. For ease of exposition, we assume training with SGD using batch size $B$ and a learning rate $\eta_{t}$. An extension to Adam is provided in Appendix C.2. The parameter update at iteration $t$ is $\bm{\theta}_{t+1}-\bm{\theta}_{t}=-\frac{\eta_{t}}{B}\sum_{j=1}^{B}\nabla_{\bm{\theta}}\ell(\bm{x}_{t}^{j};\bm{\theta}_{t})$. After the update, the logits matrix for datapoint $\bm{x}_{t}^{i}$ can be expanded using a first-order Taylor expansion:
@@
Larger Nuclear Norm indicates higher Intra-sample Diversity.
Beyond optimization utility, a larger nuclear norm can also arise from the structural diversity of token-level output distributions within a sequence. Recall Lemma 3.1: for a fixed Frobenius norm, the nuclear norm achieves its minimum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-$1$ with only one singular value, and its maximum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank with equal singular values. These two extremes correspond to distinct regimes of intra-sample diversity. The lower bound typically occurs when the model predicts only a single vocabulary token throughout the sequence (indicating repetitive generation behavior), while the upper bound is achieved when the model produces a florid and diverse vocabulary distribution across tokens (reflecting richer semantic information content). Figure 3 presents a simplified example illustrating what intra-sample diversity means in real cases. Intuitively, A larger nuclear norm implies that more vocabularies in this sequence are predicted and utilized, indicating higher prediction dispersity. Specific explanations are listed below.
Low Nuclear Norm Indicates Low rank and collinear rows (low diversity). Considering the lower bound of Lemma 3.1, since $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-1, the row vectors $\bm{l}_{n}$ are nearly collinear, i.e., $\bm{l}_{n}\approx\alpha_{n}\bm{v}$ for some fixed $\bm{v}$. This degenerate spectral structure indicates minimal diversity: the model’s predictions for different tokens are aligned along the same direction, resulting in repetitive output.
High Nuclear Norm Indicates High rank and orthogonal rows (high diversity). Considering the higher bound of Lemma 3.1, $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank and all singular values are approximately equal. This occurs when the row vectors are diverse and orthogonal with comparable norms. This flat spectrum indicates maximal diversity: each token is semantically meaningful to the output, and predictions are spread across diverse directions, capturing a wide range of semantic information.
@@ 3.2 Inter-Sample Importance Score via Low-rank Similarity Matching
Current online batch selection methods (Wang et al., 2024) only consider sample diversity within the candidate batch, which is suboptimal because the capacity of the batch is much smaller than the global datasets. To enhance global diversity, we maintain a fixed-size First-In-First-Out (FIFO) memory buffer $\bm{Q}\in\mathbb{R}^{M\times d}$, which stores representations $\bm{z}\in\mathbb{R}^{d}$ of the last $M$ samples selected for training ($M\gg B$), and measure inter-sample diversity by calculating the distance between the candidate sample and the recent training history.
@@
As mentioned, the logits matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$ is semantically meaningful for computing diversity, but directly storing the whole matrix in the memory buffer is prohibitive (e.g., $\sim 74$GB for $1024$ samples in Qwen-2.5-7B). A natural strategy is to randomly project it into low-dimensional vectors $\bm{z}_{t}^{i}\in\mathbb{R}^{d}$, while preserving pairwise distances (Johnson et al., 1984). However, a direct projection using $\bm{\Gamma}\in\mathbb{R}^{NV\times d}$ also incurs severe storage cost for the down-projection matrix ($\sim 74$GB in Qwen-2.5-7B if $d$ is set to $1024$). To avoid this, we try to factorize $\bm{\Gamma}$ into two smaller projections: $\bm{\Gamma}_{1}\in\mathbb{R}^{d_{1}\times V}$ reducing the vocabulary dimension and $\bm{\Gamma}_{2}\in\mathbb{R}^{d_{2}\times N}$ reducing the sequence length dimension, which together approximate the original projection with $d=d_{1}d_{2}$. Formally, we aim to find proper $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ that operate in the following way:
$\bm{z}=\mathrm{vec}\left(\bm{\Gamma}_{2}\cdot\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\bm{\Gamma}_{1}^{\top}\right),$ (10)
Fortunately, the subsampled randomized Fourier transform (SRFT)-style construction for $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ (see Theorem 3.2) can fulfill this role (Ailon & Chazelle, 2006; Jin et al., 2021). This approach achieves an approximate Johnson-Lindenstrauss embedding while avoiding the storage of an explicit $NV\times d$ projection matrix and reduces the computational complexity from $\mathcal{O}(NVd)$ to $\mathcal{O}\left((N+V)d\log(NV)\right)$. The complete proof is provided in Appendix C.3.
@@ 4.1 Experimental Setup
We evaluate online batch selection methods’ performance across four key domains: (1) general knowledge understanding using the MMLU (Hendrycks et al., 2021) benchmark, with auxiliary training datasets for fine-tuning and the official test set for evaluation; (2) scientific question answering using ScienceQA (Lu et al., 2022) for both fine-tuning and evaluation; (3) mathematical reasoning on GSM8K (Cobbe et al., 2021) for both fine-tuning and evaluation; and (4) code generation using CodeAlpaca-20k (Chaudhary, 2023) for training and HumanEval (Chen et al., 2021) for evaluation. All experiments are conducted using Llama-3.1-8B (Grattafiori et al., 2024) and Qwen-2.5-7B (Yang et al., 2024) as backbone models following (Zou et al., 2025).
@@
We compare UDS against the following baselines: (1) Regular, where all samples are used without data selection; (2) Random Selection, which randomly chooses samples from batches; (3) MaxLoss (Loshchilov & Hutter, 2015), which prioritizes samples with the highest training loss; (4) MaxGrad (Katharopoulos & Fleuret, 2018), which selects samples with the largest gradient norms; (5) RHO-Loss (Mindermann et al., 2022), a representative method using a reference model; and (6) GREATS (Wang et al., 2024), a state-of-the-art (SOTA) online batch selection approach. All baselines select the same fraction of data for a fair comparison.
@@
Under hardware constraints, we employ LoRA (rank=8) for SFT training. The batch size is set to $B=8$ for all datasets. The optimal data selection ratio is highly dependent on both the backbone model and the dataset; the ratios we adopt for different combinations of backbones and datasets are reported in Table 5. We report average accuracy on MMLU, ScienceQA, and GSM8K, and Pass@1 for HumanEval. We also compare training speed by throughput relative to an NVIDIA GeForce RTX 3090 GPU, following Wang et al. (2024). All benchmarks use zero-shot evaluation and are repeated four times with different random seeds. More detailed experimental settings are provided in Appendix D.
@@ Appendix A Related Work
Several studies have investigated online batch selection prior to the era of LLMs. Works like (Jiang et al., 2019; Katharopoulos & Fleuret, 2018; Loshchilov & Hutter, 2015) propose heuristic methods that select “important” samples (e.g., via high training loss, or large gradient norm). These methods are straightforward and often effective for smaller models or simpler tasks, but more recent studies (Wang et al., 2024) show that these heuristics may not scale well for LLMs. Another paradigm uses a reference model to help evaluate sample importance. For instance, (Deng et al., 2023; Mindermann et al., 2022) use a pre-trained or separately trained reference model, and select samples based on how “different” they are relative to that model. These methods tend to require extra computation (e.g. extra forward passes) and sometimes held-out data to train the reference model, which can make them costly or impractical in large-scale LLM training. (Kaddour et al., 2023) points out that these overheads can outweigh benefits in many settings. GREATS (Wang et al., 2024) is a more recent SOTA method. It selects data batches online by considering loss reduction on a validation set, and uses a “ghost inner-product” approximation (via a Taylor-expansion‐based greedy algorithm) to accelerate selection. However, the need for a validation set may be unavailable or non-representative in practical deployment (Wang et al., 2024).
@@
Most existing approaches to data selection are performed in an offline manner, where samples are filtered prior to training (Albalak et al., 2024; Xia et al., 2024; Zhou et al., 2023; Chen et al., 2023; Xie et al., 2023; Wettig et al., 2024). This is primarily due to efficiency concerns, since curating the entire dataset at every iteration is prohibitively expensive. Early approaches extend uncertainty estimation and active learning heuristics, such as filtering by entropy or difficulty, but these remain limited in capturing truly informative samples (Xu et al., 2020). A more impactful direction is large-scale pruning and deduplication: (Lee et al., 2021; Tirumala et al., 2023) show that removing near-duplicate or low-quality data improves efficiency and prevents overfitting. In instruction tuning, several works highlight that “quality outweighs quantity”, with methods like LIMA (Zhou et al., 2023), LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2022) filtering diverse, representative subsets while discarding noisy prompts. While effective to some extent, this perspective assumes that the value of each example is fixed throughout the learning process. In reality, learning is highly dynamic. Data that appear highly informative at the beginning of training may later become redundant (Wang et al., 2024). This limitation motivates serving online batch selection as a complement for dynamically assessing sample value during the training process.

### v2

@@ 1 Introduction
The rapid progress of Large Language Models (LLMs) has reshaped natural language processing and enabled impressive generalization across a wide range of domains (achiam2023gpt; brown2020language; liu2024deepseek; touvron2023llama; bai2023qwen). To further improve their performance in specialized areas, Supervised Fine-tuning (SFT) has emerged as a typical post-training paradigm. However, training on a full dataset is not always advantageous. It entails substantial computational cost and has been shown to be less effective than using a small amount of carefully curated data (albalak2024survey; xia2024less; zhou2023lima). These challenges highlight the importance of principled data selection strategies that can improve both efficiency and effectiveness in SFT.
@@
In this work, we address this challenge by focusing on online batch selection, a paradigm that dynamically evaluates sample value and performs filtering during training (wang2024greats; loshchilov2015online; katharopoulos2018not; mindermann2022prioritized). Instead of using all samples in a batch, the model assesses the importance of each sample as it is encountered and selects only a subset to participate in parameter updates. This approach enables the selection process to adapt in real time to the model’s current state and learning trajectory, improving both training efficiency and effectiveness.
@@
However, despite promising progress, current online batch selection methods still face several issues that hinder their practical deployment. These methods typically rely solely on data utility, e.g., picking sample subsets with high loss (loshchilov2015online; jiang2019accelerating) or gradient magnitude (katharopoulos2018not). While such heuristics enable capturing the critical samples to reduce the current training loss, they evaluate sample value from a limited perspective. In practice, effective selection also requires considerations of intra-sample diversity, to reflect the richness of information with fewer repetitive phrases within each training instance, and inter-sample diversity to suppress redundancy across examples by avoiding repeated training on near-duplicate content (lee2021deduplicating; tirumala2023d4). Popular literature (loshchilov2015online; katharopoulos2018not; mindermann2022prioritized) prioritizes data utility and rarely examines the selection criteria through the lens of diversity.
@@
Another primary issue involves both external dependencies and computational overhead. Some approaches (mindermann2022prioritized; wang2024greats) rely on a held-out validation set, which may be unavailable since we hardly know the exact distribution of test dataset. Other methods (mindermann2022prioritized; deng2023towards) use a reference model that is also shown to be impractical in the real world (kaddour2023no). Besides these external dependencies, several selection algorithms (katharopoulos2018not; wang2024greats; mindermann2022prioritized) introduce substantial computational costs that even exceed the expense of full-dataset training, raising efficiency concerns.
@@
Leveraging logits as basis for online scoring. To operationalize these desiderata, a key challenge lies in how to efficiently and reliably score candidate samples during training. In online batch selection, each candidate sample must be evaluated under the current model to quantify its contribution to learning. This evaluation typically requires at least a forward pass, as it can fully capture how the model presently interprets the sample and thus provide an accurate basis for selection (huang2023look; shorinwa2025survey). To achieve this efficiently (D3) without external resources (D2), we must avoid expensive gradient computations for each candidate sample and leverage intrinsic signals already available during forward propagation. These considerations collectively justify leveraging the model’s output logits, which are naturally produced during the forward pass and encode rich information about both sample utility and diversity (qiu2024semantic; geng2023survey). Specifically, we compute an intra-sample importance score based on the nuclear norm of the logits, which captures the utility and intra-sample diversity of a sample, and an inter-sample importance score using low-dimensional similarity matching against previously selected samples to quantify inter-sample diversity (D1). Integrating these scores allows the training process to balance exploitation of high-utility samples with exploration of less-visited regions in the data distribution, thereby improving the overall effectiveness of online batch selection.
@@ 3.1 Intra-sample Importance Score via Nuclear Norm
(horn2012matrix) For any matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$, the following inequality holds:
@@
Larger Nuclear Norm Positively Correlates With Higher Optimization Utility.
Traditional notions of sample difficulty—such as maximum loss (loshchilov2015online) or maximum gradient (katharopoulos2018not)—often misalign with actual training dynamics (wang2024greats). Instead, we characterize sample importance through its potential contribution to loss reduction, termed its optimization utility. For ease of exposition, we assume training with SGD using batch size $B$ and a learning rate $\eta_{t}$. An extension to Adam is provided in Appendix C.2. The parameter update at iteration $t$ is $\bm{\theta}_{t+1}-\bm{\theta}_{t}=-\frac{\eta_{t}}{B}\sum_{j=1}^{B}\nabla_{\bm{\theta}}\ell(\bm{x}_{t}^{j};\bm{\theta}_{t})$. After the update, the logits matrix for datapoint $\bm{x}_{t}^{i}$ can be expanded using a first-order Taylor expansion:
@@
Figure 3: Strong correlation between $rank(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t}))$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ during the training process. Each point represents the correlation coefficient calculated from $B$ pairs of values within a single batch. We extract the result using Qwen-2.5-7B on MMLU.
@@
Larger Nuclear Norm Positively Correlates With Higher Intra-sample Diversity.
Beyond optimization utility, a larger nuclear norm can also arise from the structural diversity of token-level output distributions within a sequence. Recall Lemma 3.1: for a fixed Frobenius norm, the nuclear norm achieves its minimum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-$1$ with only one singular value, and its maximum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank with equal singular values. These two extremes correspond to distinct regimes of intra-sample diversity. The lower bound typically occurs when the model predicts only a single vocabulary token throughout the sequence (indicating repetitive generation behavior), while the upper bound is achieved when the model produces a florid and diverse vocabulary distribution across tokens (reflecting richer semantic information content). Figure 4 presents a simplified example illustrating what intra-sample diversity means in real cases. Intuitively, A larger nuclear norm implies that more vocabularies in this sequence are predicted and utilized, indicating higher prediction dispersity. We additionally provide an empirically strong correlation between $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ and $rank(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t}))$ in Figure 3 to support this. Specific explanations are listed below.
Low Nuclear Norm Positively Correlates with Low rank and collinear rows (low diversity). Considering the lower bound of Lemma 3.1, since $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-1, the row vectors $\bm{l}_{n}$ are nearly collinear, i.e., $\bm{l}_{n}\approx\alpha_{n}\bm{v}$ for some fixed $\bm{v}$. This degenerate spectral structure indicates minimal diversity: the model’s predictions for different tokens are aligned along the same direction, resulting in repetitive output.
High Nuclear Norm Positively Correlates with High rank and orthogonal rows (high diversity). Considering the higher bound of Lemma 3.1, $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank and all singular values are approximately equal. This occurs when the row vectors are diverse and orthogonal with comparable norms. This flat spectrum indicates maximal diversity: each token is semantically meaningful to the output, and predictions are spread across diverse directions, capturing a wide range of semantic information.
@@ 3.2 Inter-Sample Importance Score via Low-rank Similarity Matching
Current online batch selection methods (wang2024greats) only consider sample diversity within the candidate batch, which is suboptimal because the capacity of the batch is much smaller than the global datasets. To enhance global diversity, we maintain a fixed-size First-In-First-Out (FIFO) memory buffer $\bm{Q}\in\mathbb{R}^{M\times d}$, which stores representations $\bm{z}_{t}^{i}\in\mathbb{R}^{d}$ of the last $M$ samples selected for training ($M\gg B$), and measure inter-sample diversity by calculating the distance between the candidate sample and the recent training history.
@@
As mentioned, the logits matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$ is semantically meaningful for computing diversity, but directly storing the whole matrix in the memory buffer is prohibitive (e.g., $\sim 74$GB for $1024$ samples in Qwen-2.5-7B). A natural strategy is to randomly project it into low-dimensional vectors $\bm{z}_{t}^{i}\in\mathbb{R}^{d}$, while preserving pairwise distances (johnson1984extensions). However, a direct projection using $\bm{\Gamma}\in\mathbb{R}^{NV\times d}$ also incurs severe storage cost for the down-projection matrix ($\sim 74$GB in Qwen-2.5-7B if $d$ is set to $1024$). To avoid this, we try to factorize $\bm{\Gamma}$ into two smaller projections: $\bm{\Gamma}_{1}\in\mathbb{R}^{d_{1}\times V}$ reducing the vocabulary dimension and $\bm{\Gamma}_{2}\in\mathbb{R}^{d_{2}\times N}$ reducing the sequence length dimension, which together approximate the original projection with $d=d_{1}d_{2}$. Formally, we aim to find proper $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ that operate in the following way:
$\bm{z}_{t}^{i}=\mathrm{vec}\left(\bm{\Gamma}_{2}\cdot\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\bm{\Gamma}_{1}^{\top}\right),$ (10)
Fortunately, the subsampled randomized Fourier transform (SRFT)-style construction for $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ (see Theorem 3.2) can fulfill this role (ailon2006approximate; jin2021faster). This approach achieves an approximate Johnson-Lindenstrauss embedding while avoiding the storage of an explicit $NV\times d$ projection matrix and reduces the computational complexity from $\mathcal{O}(NVd)$ to $\mathcal{O}\left((N+V)d\log(NV)\right)$. The complete proof is provided in Appendix C.3.
@@ 4.1 Experimental Setup
We evaluate online batch selection methods’ performance across four key domains: (1) general knowledge understanding using the MMLU (hendryckstest2021) benchmark, with auxiliary training datasets for fine-tuning and the official test set for evaluation; (2) scientific question answering using ScienceQA (lu2022learn) for both fine-tuning and evaluation; (3) mathematical reasoning on GSM8K (cobbe2021gsm8k) for both fine-tuning and evaluation; and (4) code generation using CodeAlpaca-20k (codealpaca) for training and HumanEval (chen2021evaluating) for evaluation. All experiments are conducted using Llama-3.1-8B (grattafiori2024llama) and Qwen-2.5-7B (yang2024qwen2) as backbone models following zou2025flylora.
@@
We compare UDS against the following baselines: (1) Regular, where all samples are used without data selection; (2) Random Selection, which randomly chooses samples from batches; (3) MaxLoss (loshchilov2015online), which prioritizes samples with the highest training loss; (4) MaxGrad (katharopoulos2018not), which selects samples with the largest gradient norms; (5) RHO-Loss (mindermann2022prioritized), a representative method using a reference model; and (6) GREATS (wang2024greats), a state-of-the-art (SOTA) online batch selection approach. All baselines select the same fraction of data for a fair comparison.
@@
Under hardware constraints, we employ LoRA (rank=8) for SFT training. The batch size is set to $B=8$ for all datasets. The optimal data selection ratio $\alpha$ is highly dependent on both the backbone model and the dataset; the ratios we adopt for different combinations of backbones and datasets are reported in Table 5. The default choice of the buffer size $M=1024$, the down-projection dimension is $d_{1}=128$, $d_{2}=8$, and we use this across all the experiments. For detailed sensitivity analysis, please refer to Appendix B.1. We report average accuracy on MMLU, ScienceQA, and GSM8K, and Pass@1 for HumanEval. We also compare training speed by throughput relative to an NVIDIA GeForce RTX 3090 GPU, following wang2024greats. All benchmarks use zero-shot evaluation and are repeated four times with different random seeds. More detailed experimental settings are provided in Appendix D.
@@ Appendix A Related Work
Several studies have investigated online batch selection prior to the era of LLMs. Works like (jiang2019accelerating; katharopoulos2018not; loshchilov2015online) propose heuristic methods that select “important” samples (e.g., via high training loss, or large gradient norm). These methods are straightforward and often effective for smaller models or simpler tasks, but more recent studies (wang2024greats) show that these heuristics may not scale well for LLMs. Another paradigm uses a reference model to help evaluate sample importance. For instance, (deng2023towards; mindermann2022prioritized) use a pre-trained or separately trained reference model, and select samples based on how “different” they are relative to that model. These methods tend to require extra computation (e.g. extra forward passes) and sometimes held-out data to train the reference model, which can make them costly or impractical in large-scale LLM training. (kaddour2023no) points out that these overheads can outweigh benefits in many settings. GREATS (wang2024greats) is a more recent SOTA method. It selects data batches online by considering loss reduction on a validation set, and uses a “ghost inner-product” approximation (via a Taylor-expansion‐based greedy algorithm) to accelerate selection. However, the need for a validation set may be unavailable or non-representative in practical deployment (wang2024greats).
@@
Most existing approaches to data selection are performed in an offline manner, where samples are filtered prior to training (albalak2024survey; xia2024less; zhou2023lima; chen2023alpagasus; xie2023data; wettig2024qurating). This is primarily due to efficiency concerns, since curating the entire dataset at every iteration is prohibitively expensive. Early approaches extend uncertainty estimation and active learning heuristics, such as filtering by entropy or difficulty, but these remain limited in capturing truly informative samples (xu2020curriculum). A more impactful direction is large-scale pruning and deduplication: (lee2021deduplicating; tirumala2023d4) show that removing near-duplicate or low-quality data improves efficiency and prevents overfitting. In instruction tuning, several works highlight that “quality outweighs quantity”, with methods like LIMA (zhou2023lima), LESS (xia2024less), and Self-Instruct (wang2022self) filtering diverse, representative subsets while discarding noisy prompts. Recent advances further explore offline data selection through various sampling strategies: sachdeva2024train introduce LLM-as-judge method and density-based sampling; axiotis2024data propose clustering-based sensitivity sampling; and deb2025fishersft leverage information gain estimation to identify high-value examples. While effective to some extent, this perspective assumes that the value of each example is fixed throughout the learning process. In reality, learning is highly dynamic. Data that appear highly informative at the beginning of training may later become redundant (wang2024greats). This limitation motivates serving online batch selection as a complement for dynamically assessing sample value during the training process.